40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Sep 9 11:59:58 2011, kiwamu, Update, IOO, PZT1 doesn't work 
    Reply  Fri Sep 9 13:29:09 2011, kiwamu, Update, IOO, a history of PZT mirror 
Message ID: 5368     Entry time: Fri Sep 9 11:59:58 2011     Reply to this: 5369
Author: kiwamu 
Type: Update 
Category: IOO 
Subject: PZT1 doesn't work 

Last night I noticed that PZT1 didn't work properly

I am not sure what is going on. Today I will try localizing the cause of the problem.

As far as I remember it was perfectly working at the time just after we readjusted the OSEMs on MC1 and MC3 (Aug 23th)

 

The symptoms are :

  + No response to both pitch and yaw control from EPICS (i.e. C1:LSC-PZT1_X and C1:LSC-PZT1_Y)

  + When a big value (-3 or so) from EPICS was applied, the PZT1 mirror suddenly jumped.

     However it turned out it just corresponded to a state where OOR (Out Of Range) LED lights up.

 

I did some brief checks :

  + checked the voltage going into the HV amplifiers' "MOD" input. Those are the voltage coming out from DACs and controlled from EPICS.

   --> looked healthy. They went from -10 to 10 V as expected (although the HV amp takes up to only +/-5V).

  + swapped the ''MOD" input cables such that C1:LSC-PZT1 controls the PZT2 HV and vice versa.

    --> The PZT2 mirror was still controlable, but the PZT1 mirror still didn't move. So the DAC and EPICS are innocent.

  + swapped the D-dub cables, which are directly going into the feedthroughs, such that the PZT1 HV drives the PZT2 mirrors and vice versa.

    --> the PZT2 mirror became unable to be controlled. For the PZT1 mirror, only PITCH worked smoothly.


ELOG V3.1.3-